OpenStudy (anonymous):

Is the equation true, false, or open? Explain. 9p + 8 = 10p + 7

OpenStudy (tkhunny):

Explain by solving. 9p + 8 = 10p + 7 Subtract 9p 8 = p + 7 Subtract 7 1 = p Okay, there IS a solution, but it is nto ALWAYS true. It is true ONLY if p = 1. This makes the original equation Conditionally True.

OpenStudy (anonymous):

Thank you. I am stumped with these assignments can you help me for a second? Please?..

OpenStudy (tkhunny):

How do you know I'm not afraid of English Teachers?

OpenStudy (anonymous):

Not English Teachers... Englsh Teachers.

OpenStudy (anonymous):

Well I'm no English Teacher, yet. But I am aspiring to become one. I just need help with math.

OpenStudy (anonymous):

and Geoffb, I know how to spell; the system wouldn't let me have too many letters. Worry about what's on your account, and helping the other students. Thanks. Have a blessed night!

OpenStudy (anonymous):

Oh, settle down. It's a joke, obviously.

OpenStudy (tkhunny):

I see. Then "Conditionally True" is no good. The only choice from that list is "Open Sentence"